OpenStudy (anonymous):

Which of the graphs represents the inequality y – 2x < 3?

OpenStudy (atlas):

DO you know how to draw a graph for the line y -2x=3

OpenStudy (anonymous):

I dont see how too get the answer though?? Its diving me crazy cause i have tried everythang i no to do.

OpenStudy (atlas):

i asked you can you draw the line y-2x=3 in a graph?

OpenStudy (anonymous):

no?

OpenStudy (atlas):

ok......i will tell you a simple way of doing so

OpenStudy (anonymous):

Okay.

OpenStudy (atlas):

take any value of x. lets say x=1 what is the value of y now

OpenStudy (anonymous):

1

OpenStudy (atlas):

Nope...check again

OpenStudy (anonymous):

-2 i think im sorry i suck at this..

OpenStudy (atlas):

y -2x =3 I put x=1 so, y -2(1) =3 or y -2 =3 so y =??

OpenStudy (atlas):

no you are good......just think a bit

OpenStudy (atlas):

you know when two sides are equal you can add and subtract equal numbers from it and they still will remain equal. so if y-2 =3 I can add 2 on both sides and they still remain equal

OpenStudy (atlas):

Can you get y now?

OpenStudy (anonymous):

So it would stay at 2 so then it would be y=2+5 ?

OpenStudy (atlas):

no no no

OpenStudy (anonymous):

:(

OpenStudy (atlas):

y-2 =3 so i add 2 on both sides y-2+2 =3+2 y=5

OpenStudy (anonymous):

Oh i get it cause of corse the 2 would cancel out and then it would be left at 3+2 which is 5. so then hoe would you be able to graph it?

OpenStudy (atlas):

Now since x=1 and y=5 satisfy the equation (1,5) is a point on the graph

OpenStudy (anonymous):

* sorry i ment how?

OpenStudy (anonymous):

Yes. But there has too be another point to correct?

OpenStudy (atlas):

Yes exactly! so i take another value of x. lets say x=2 and find the value of y again

OpenStudy (anonymous):

So then it would be y=-2x=2 +2 +2 y=-1x-2?

OpenStudy (atlas):

nope we do it this way: y-2x =3 I take x =2 so, y - 2*2 =3 or y-4=3 Now can you find y??

OpenStudy (anonymous):

y=-1

OpenStudy (atlas):

Nope again y -4 =3 Now as I told you before when two sides are equal I can add or subtract any numbers from both sides and they will still remain equal. So I add 4 on both sides y-4+4 = 3+4 y =7

OpenStudy (anonymous):

That would be rite i aded the wrong side together. Thats were i was wrong. I add the oppisite side togeter and thats how im gettin the wrong answer.

OpenStudy (atlas):

You just have to work on your calculations a bit :D

OpenStudy (anonymous):

Well thanks so much dollface yo was a BIG help as you can see i was lost!!

OpenStudy (atlas):

Ok so we have two points that we know satisfy the equation (1,5) and (2,7) Can you now find a line on the graph that passes through these points?

OpenStudy (anonymous):

Yes.

OpenStudy (atlas):

K now this line is y-2x =3

OpenStudy (atlas):

and this line divides the whole plane into two parts

OpenStudy (atlas):

One part is y-2x >3 and the other part is y-2x<3

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Yes no i have the correct answer and finally passed the quiz!

OpenStudy (atlas):

Oh great!

OpenStudy (anonymous):

Thanks so much!!!!

OpenStudy (atlas):

You are welcome :)

OpenStudy (anonymous):

:)